OpenStudy (anonymous):

please help me with just answers please :) Question 5 Which is the equation to the line passing through points (2, –7) and (5, –1)? Answer A) y = – 2x + 6 B) y = 2x – 11 C) y = 2x – 8 D) y = – 2x + 9 Question 6 What is the equation of a vertical line passing through (1, 7)? Answer A) y = 1 B) x = 1 C) x = 7 D) y = 7 Question 7 What is the equation of the line that passes through (12, 4) and is perpendicular to the graph of y = –x – 2? a) y = –x + 20 b) y = x – 12 c) y = x – 5 d) y = –x + 13

OpenStudy (sasogeek):

why would u want just the answers... do u have a method to solving them already?

OpenStudy (anonymous):

because im just looking for the answer..

OpenStudy (sasogeek):

what if i give u the wrong answers and tell you, hey, those are the answers... would u walk away thinking u have the right answers? don't u think u should know how I get my answers?

OpenStudy (anonymous):

\[y=2x-11\] for the first one

OpenStudy (anonymous):

thanks for the first one :)

OpenStudy (amistre64):

trust me, just the answers is about it :)

OpenStudy (amistre64):

y=7 for the next i think

OpenStudy (anonymous):

\[x=1\] for the next one

OpenStudy (amistre64):

vertical vs horizontal, i always gotta think about that one

OpenStudy (sasogeek):

what are u guys doing, amistre says y=7 is the answer, satellite says x=1 is the answer... who's to believe?

OpenStudy (amistre64):

id go with the smart guy ;)

OpenStudy (anonymous):

horizontal... like horizon 

OpenStudy (anonymous):

wait #6 you think its 1 or 7?

OpenStudy (anonymous):

\[x=1\]

OpenStudy (amistre64):

i remember that picture

OpenStudy (sasogeek):

i think it's both

OpenStudy (anonymous):

it cant be both answers..

OpenStudy (sasogeek):

:P you can go with my answer if u want :)

OpenStudy (sasogeek):

so then don't answer it, u wanted an answer, i gave u one :)

OpenStudy (anonymous):

and what is your answer

OpenStudy (anonymous):

let me repeat \[x=1\]

OpenStudy (anonymous):

ok thanks so much :)

OpenStudy (amistre64):

we can rule out y=7 becasue its simply not sunny enough

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok #7?..

OpenStudy (anonymous):

and last one is not one of your choices perpendicular to \[y=x-2\] thought \[(12,4)\] the line would be \[y-4=x-12\] or \[y=x-8\] so either there is a typo or the question is incorrect

OpenStudy (anonymous):

answer to "perpendicular to y = -x -2 though (12,4) is \[y=x-8\]

OpenStudy (anonymous):

ok thanks so much

OpenStudy (anonymous):

yw